OpenStudy (anonymous):

hoose the correct simplification of (6x^4y^2z^3)(3x^4y^3z^5). Answer 18x^16y^6z^15 18x^8y^5z^8 9x^8y^5z^8 9x^16y^6z^15

OpenStudy (anonymous):

First, multiply the constant in the front: 6*3 = 18. So, the last two are out. Then, multiply like terms (that have the same base) and add the exponents: x^4 x^4 = x^8, y^3 y^2 = y^5 and z^5 z^3 = z^8. So, the second choice is the correct one.